OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?

OpenStudy (anonymous):

3x - 2y = -6 2y - 6 = 3x 2y = 3x + 6 y = (3/2)x + 3 So, the y-intercept is 3.

OpenStudy (luigi0210):

Get your equation in slope-intercept form, first.

OpenStudy (amistre64):

parallel lines only differ at most by a constant

OpenStudy (amistre64):

the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6 is: 3x – 2y = k; such that k is determined by the stated point